Two dice are rolled. Find the probability of getting a 5 on one die or on both dice.
Answer by greenestamps(13208)



The opposite (or complement) of getting a 5 on one die or on both dice is not getting a 5 on either one. So

P(5 on one die or both dice) = 1 - P(both dice not 5) = 1 - (5/6)(5/6) = 1-25/36 = 11/26.

OR....

P(5 on one die or both dice) = P(5 on first die) + P(5 on second die) - P(5 on both dice) = 1/6 + 1/6 - 1/36 = 11/36.